# Evaluating GPT-V4 (GPT-4 with Vision) on Detection of Radiological Findings on Chest Radiographs

### v2

@@ 3.3 Inter-rater Agreement
The study compared each reader’s outcomes to the reference standard to evaluate their interpretations (Table 1). On the NIH dataset, the IRA for Reviewer A was 0.78, for Reviewer C was -0.06, and for Reviewer D was 0.96. On the MIDRC dataset, the IRA for Reviewer B was 0.96, for Reviewer C was 0.82, and for Reviewer D was 0.99.
Table 1: Reviewer Agreement (Cohen’s Kappa) for NIH and MIDRC. The Inter-rater Agreement (IRA) between each reviewer and the reference standard table for both NIH and MIDRC datasets. $k=1$ indicates perfect agreement among reviewers and the reference standard; $k=0$ indicates that agreement is no better than chance; $k=\ -1$ indicates perfect disagreement.
NIH Results MIDRC Results
Reviewer A 0.78 NA
Reviewer B NA 0.96
Reviewer C -0.06 0.82
Reviewer D 0.96 0.99
@@ 4 Discussion
The emergence of multimodal large language models (LLMs) that can understand both text and images, such as GPT-4 with Vision (GPT-4V), shows potential for automated image-text pair generation. However, applying these models to real-world data is yet to be thoroughly examined. This study assessed the feasibility of using GPT-4V to detect radiological findings from chest radiographs in both zero-shot and few-shot learning contexts. The results (average PPV of 5.53/45 (12.3%) and average TPR of 2.60/45 (5.8%)) demonstrated that radiological findings tables generated by the GPT-4V model still need to be prepared for use in clinical practice. We acknowledged a limitation in employing GPT-4 for converting radiological reports into a structured table, where inaccuracies in ICD-10 code assignments and distinguishing between radiological findings and conclusions may impact the data’s reliability and interpretability. A notable limitation of the GPT-4V output was its failure to detect several radiological conditions based on corresponding codes in the International Classification of Diseases, Tenth Revision (ICD-10). Overall, the top three findings that GPT-4V could not detect were “endotracheal tube”, “central venous catheter”, and “degenerative changes of osseous structures”. Conversely, GPT-4V most accurately detected findings such as “chest drain”, “air-space disease”, and “lung opacity”.
@@
Our study had limitations. First, the few-shot training is that it may be more prone to generating ICD-10 codes already in the provided examples, potentially reducing the diversity of ICD-10 codes generated. Second, we were unable to access other multimodal LLMs that support image inputs. Consequently, we lacked comparative data with respect to the results of GPT-4v. Finally, due to the limited size of the dataset and relatively low level of inter-rater agreement among the radiologists, analysis by GPT-4V may have been challenging.
In conclusion, GPT-4 with Vision (GPT-4V) showed promise in understanding natural images but had limited effectiveness in interpreting real-world chest radiographs. Our results highlight the need for additional comprehensive development and assessments prior to incorporating the GPT-4V model into clinical practice routines. Task-specific, fine-tuned multimodal LLMs or foundation models are urgently needed for this purpose, although it is not necessarily the best solution. To yield robust and generalizable results, we plan to explore larger and more diverse datasets using real-world data in future studies. This will involve including multiple modalities, such as brain CT scans and MRIs, to conduct a more thorough evaluation of GPT-4V’s performance.

### v3

@@ 3.3 Inter-rater Agreement
The study compared each reader’s outcomes to the reference standard to evaluate their interpretations (Table 1). On the NIH data set, the interrater agreement was 0.78 for reader A, -0.06 for reader C, 0.96 for reader D. On the MIDRC data set, the interrater agreement was 0.96 for reader B, 0.82 for reader C, and 0.99 for reader D.
Table 1: Reader Agreement for NIH and MIDRC. Note. —Interrater Agreement between each reader and the reference standard table for both NIH and MIDRC data sets were accessed with Cohen $\kappa$ statistics. $\kappa$ values were interpreted as follows: $\kappa=1$ indicates perfect agreement among reviewers and the reference standard; $\kappa=0$ indicates that agreement is no better than chance; $\kappa=\ -1$ indicates perfect disagreement.MIDRC = Medical Imaging and Data Resource Center, NA = not applicable, NIH = National Institutes of Health.
NIH $\kappa$ Score MIDRC $\kappa$ Score
Reader A 0.78 NA
Reader B NA 0.96
Reader C -0.06 0.82
Reader D 0.96 0.99
@@ 4 Discussion
The emergence of multimodal large language models (LLMs) that can understand both text and images, such as OpenAI’s GPT-4V [1], shows potential for automated image-text pair generation. However, applying these models to real-world data is yet to be thoroughly examined. This study assessed the feasibility of using GPT-4V to detect radiologic findings from chest radiographs in both zero-shot and few-shot learning contexts. The results (average PPV, 5.53 of 45 [12.3%]; average TPR, 2.60 of 45 [5.8%]) demonstrated that radiologic findings tables generated by GPT-4V still need further preparation for use in clinical practice. We acknowledge a limitation in employing GPT-4 for converting radiologic reports into a structured table, where inaccuracies in the International Statistical Classification of Diseases, Tenth Revision (ICD-10) code assignments and distinguishing between radiologic findings and conclusions may impact the reliability and interpretability of the data. A notable limitation of the GPT-4V output was its failure to detect several clinical conditions based on corresponding ICD-10 codes. Overall, the top three findings that GPT-4V could not detect were “endotracheal tube,” “central venous catheter,” and “degenerative changes of osseous structures.” Conversely, GPT-4V most accurately detected findings such as “chest drain,” “air-space disease,” and “lung opacity.”
@@
Our study had limitations. First, few-shot training is that it may be more prone to generating ICD-10 codes already in the provided examples, potentially reducing the diversity of ICD-10 codes generated. Second, we were unable to access other multimodal LLMs that support image inputs. Consequently, we lacked comparative data with respect to the results of GPT-4v. Finally, due to the limited size of the data set and relatively low level of inter-rater agreement among the radiologists, analysis by GPT-4V may have been challenging.
In conclusion, GPT-4V has shown promise in understanding natural images but had limited effectiveness in interpreting real-world chest radiographs. Our results highlight the need for additional comprehensive development and assessments prior to incorporating the GPT-4V model into clinical practice routines. Task-specific, fine-tuned multimodal LLMs or foundation models are urgently needed for this purpose, although it is not necessarily the best solution. To yield robust and generalizable results, we plan to explore larger and more diverse data sets using real-world data in future studies. This will involve including multiple modalities, such as brain CT scans and MRIs, to conduct a more thorough evaluation of the performance of GPT-4V.
@@ Acknowledgment
Appendix A Prompt to convert the free-text report into a table of radiological findings
Read the report and look for any important clinical findings. Provide a summary in a table format where the positive clinical conditions are 1 and the negative clinical conditions are 0. Designate each condition as left side, right side, or bilateral, Provide an ICD-10 code in a separate column for positive findings only or N/A if not applicable.
Table columns include: [Exam No., Finding No., Clinical Finding, Left Side. Right Side, Bilateral, Midline, lCD-10 Code, iCD-10 Description]
Additional instructions:
1. Normal findings should be excluded from each table
2. Group similar findings together where possible for each table
3. Create a table
Appendix B Reference standard radiological findings
Supplementary Table 1: Reference standard radiological findings
Radiologic Finding Location ICD-10 ICD-10 Description
Endotracheal tube placement Midline Z93.0 Tracheostomy status
Pulmonary opacities Bilateral J84.9 Interstitial pulmonary disease, unspecified
Pleural effusion C Left Side J90 Pleural effusion, not elsewhere classified
Pneumonia Bilateral J18.9 Pneumonia, unspecified organism